OpenStudy (anonymous):

Give an example of a rational function that has a horizontal asymptote of y = 2/9?

terenzreignz (terenzreignz):

If you want a horizontal asymptote of y = c (where c is any constant) a worthy candidate is \[\Large f(x) = \frac1{x - c}\]

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

can i use the example of y=(2x+5) / 9x

OpenStudy (anonymous):

if you wanted a "vertical" asymptote that would do, but for a horizontal asymptote you want yes

terenzreignz (terenzreignz):

Crud... sorry guys, fail :3

OpenStudy (anonymous):

\[y=\frac{2x+5}{9x}\] will do

OpenStudy (anonymous):

how would i go to find the asymptote of this

terenzreignz (terenzreignz):

More like \[\Large f(x) = \frac1{x}+c\]

OpenStudy (anonymous):

do i set both the denominator and numerator to zero?

OpenStudy (anonymous):

numerator and denominator have the same degree and the ratio of the leading coefficients is \(\frac{2}{9}\)

terenzreignz (terenzreignz):

But the asymptote was given in the question?

OpenStudy (anonymous):

yes but i need to find an equation that fits the asymptote

OpenStudy (anonymous):

for example, the horizontal asymptotes of \[y=\frac{2x^2+6x-8}{9x^2+7x-2}\] is also \(\frac{2}{9}\)

OpenStudy (anonymous):

you found one already

OpenStudy (anonymous):

how do i find the horizontal asymptote?

terenzreignz (terenzreignz):

It's \(\large y = \frac29\) as it has been written?

OpenStudy (anonymous):

i know but step by step

terenzreignz (terenzreignz):

Ahh, I see, well, if you the equation you gave... \[y=\frac{2x+5}{9x}\] Then just take the limit as x gets bigger without bound... \[\large \lim_{x\rightarrow\infty}\frac{2x+5}{9x}\] And you'll find the (horizontal) asymptote...

OpenStudy (anonymous):

y = (2x + 5)/(9x) lim y = 2/9 x-> ∞ and lim y = 2/9 x-> -∞ but y never equals 2/9

OpenStudy (anonymous):

like this?

terenzreignz (terenzreignz):

Yes, like that.

OpenStudy (anonymous):

ok thank you very much :)

terenzreignz (terenzreignz):

Sorry about that first one... gotta think these things through :3 -------------------------------- Terence out